### v1

# How do Pre-Trained Models Support Software Engineering? An Empirical Study in Hugging Face

Abstract
Open-Source Pre-Trained Models (PTMs) provide extensive resources for various Machine Learning (ML) tasks, yet these resources lack a classification tailored to Software Engineering (SE) needs. To address this gap, we derive a taxonomy encompassing 147 SE tasks and apply an SE-oriented classification to PTMs in a popular open-source ML repository, Hugging Face (HF). Our repository mining study began with a systematically gathered database of PTMs from the HF API, considering their model card descriptions and metadata, and the abstract of the associated arXiv papers. We confirmed SE relevance through multiple filtering steps: detecting outliers, identifying near-identical PTMs, and the use of Gemini 2.0 Flash, which was validated with five pilot studies involving three human annotators. This approach uncovered 2,205 SE PTMs. We find that code generation is the most common SE task among PTMs, primarily focusing on software implementation, while requirements engineering and software design activities receive limited attention. In terms of ML tasks, text generation dominates within SE PTMs. Notably, the number of SE PTMs has increased markedly since 2023 Q2. Our classification provides a solid foundation for future automated SE scenarios, such as the sampling and selection of suitable PTMs.
Index Terms:
Software Engineering Pre-trained Models, Hugging Face, Model classification
I Introduction
The exponential growth of open-source platforms such as Hugging Face (HF) [1] has enhanced access to Machine Learning (ML) assets [2]. In particular, HF is distinguished by its extensive and fast-growing collection of Pre-Trained Models (PTMs)[3, 4], compared to other platforms [5][6]. In 2024 alone, HF saw an average of 2,199 new PTMs published per day and over six million daily downloads, based on our analysis of data retrieved through the platform’s public API. Moreover, it is estimated that a new repository is created every 15 seconds [7], reflecting trends observed in prior studies that reported rapid increases in repository creation and model contributions [8]. Despite this growth, the current categorization of HF resources remains largely general-purpose, often targeting ML domains as broad as computer vision or natural language processing. This generic taxonomy fails to address the specific and diverse needs of the Software Engineering (SE) community, whose requirements differ significantly from those of other fields [9]. Importantly, SE goes beyond mere coding, encompassing a set of activities structured around the well-known Software Development Life Cycle (SDLC) [10, 11]. This cycle comprises five main activities: requirements engineering, software design, software implementation, software quality assurance, and software maintenance, thereby covering the entire process from initial specification to long-term system evolution.
Other platforms like Papers with Code [12] also provide structured connections between research papers and code implementations, offering valuable categorizations of ML tasks. However, these are primarily oriented toward traditional ML domains. Even though the platform includes an SE-related category, it combines general ML tasks (e.g., reinforcement learning, semantic segmentation) with SE-specific tasks (e.g., code generation, program repair), limiting its utility for SE-focused research and practice.
The absence of an SE-specific categorization in open-source model repositories [13] presents a significant barrier. Not only limits the efficient application of ML in SE, but it also hinders the continuous integration of new PTMs into a structured framework. This gap has several consequences for Automated Software Engineering (ASE) practitioners and researchers. In the absence of tailored filtering mechanisms, practitioners must manually sift through thousands of PTMs to sample those relevant for their tasks, an effort that is both time-consuming and error-prone, ultimately slowing progress and reducing reproducibility [14]. Moreover, without a task-aligned taxonomy, it becomes difficult to establish consistent benchmarks, which are essential for the advancement of automated development pipelines. More broadly, this lack of structured cataloguing prevents the SE community from systematically identifying relevant resources, evaluating progress, and enabling full automation in the multi-agent environment.
Given the dynamic and evolving nature of both ML and SE [15], there is a clear need for a flexible and domain-specific classification system. Such a system would support the effective sampling, selection, and integration of PTMs tailored to the full SDLC. Prior studies have taken important steps toward organizing ML resources for SE, but each presents limitations that leave a significant gap unaddressed. Hou et al. [16] proposed a taxonomy by categorizing LLM-related papers into SE tasks and SDLC activities through a systematic literature review. However, their approach is grounded in static academic sources and lacks connection to actual open-source repositories like HF. Di Sipio et al. [13] extracted SE-relevant information from HF model cards using a semi-automated method. Yet, their validation was limited to only three PTMs, and their taxonomy does not align with the SDLC, reducing its applicability. Similarly, Yang et al. [17] curated code-focused LLMs from HF, providing useful insights into that specific subset, but their scope excludes broader SE tasks and lacks an SDLC perspective. Wang et al. [18] surveyed benchmarks for code LLMs and agents around the SDLC. However, their work does not address the classification of other types of PTMs.
In this work, we address these limitations by introducing an automated classification framework for SE tasks across the SDLC. Our approach systematically maps PTMs from HF to a taxonomy of 147 SE tasks, built on prior research and established guidelines. We analyze multiple sources of descriptive content, including model card documentation, metadata, and abstracts from associated arXiv papers. The classification pipeline begins with the extraction of candidate task tokens from textual descriptions and applies several refinement steps to improve precision. These include outlier detection to remove inconsistent entries, similarity-based deduplication to eliminate near-identical variants, and an LLM support to ensure that the assigned activity reflects the model’s described purpose.
Our main contributions are as follows.
  • (a)
    An automated classification framework that accesses the HF API, processes, filters, and classifies PTMs according to the full SDLC.
  • (b)
    An extensive analysis of SE PTMs, covering their evolution, task distribution, model reuse, and reported attributes.
  • (c)
    A validated, publicly available dataset of 2,205 SE PTMs.
Our results offer several insights of particular relevance to the ASE community. While the number of SE PTMs in HF has grown significantly in the past few years, we find that they are heavily concentrated on the mid to late stages of the SDLC. By providing a reproducible, SDLC-aligned classification of PTMs, our work supports ASE researchers and practitioners by automating the identification and selection of the most appropriate PTMs for specific tasks. This automation reduces manual effort, improves consistency in benchmarking, and accelerates the adoption of ML-driven tools within ASE workflows, ultimately fostering more effective and targeted automation across the SDLC
This article is structured as follows. Section II reviews related work on PTMs and their application in SE; Section III describes the methodology used to classify PTMs hosted on HF; Section IV presents the main results and analysis; Section V discusses the findings and explores implications; Section VI identifies threats to validity; and Section VII concludes the article and suggests future research directions.
TABLE I: Summary of relevant aspects in related work.
Study Approach [19] Data Source SDLC Auto Classification Information Used Initial Resources Identified Resources
[16] Systematic Review Published Literature 88 SE tasks across 6 SE activities Paper content 218,765 papers 395 papers
[13] Repository Mining HF dump 6 macro-tasks Model card 381,240 PTMs Tested on BERT, RoBERTa, and T5
[17] Repository Mining HF API 9 categories Model documentation 1,566 PTMs 366 PTMs
[18] Systematic Review Published Literature 30 SE tasks across 6 SE activities Academic publications 1,347 papers 181 benchmarks from 461 relevant papers
This study Repository Mining HF API + arXiv 147 SE tasks across 5 SE activities Model card description, metadata, and arXiv abstract 1,533,973 PTMs 2,205 PTMs
  • SDLC indicates alignment with the Software Development Life Cycle; Auto indicates whether the work is fully automated.
II Related Work
Key aspects of the related work are summarized in Table I. A systematic literature review conducted by Hou et al. [16] analyzed 395 research papers from January 2017 to January 2024 and categorized Large Language Models (LLMs) into SE tasks. These tasks were grouped into SE activities according to the SDLC. The study identified 88 distinct SE tasks, revealing that LLMs are most commonly applied in the software development phase, with code generation and program repair being the most prevalent tasks for employing LLMs in software development and maintenance activities. In contrast, the earlier SDLC phases, such as requirements engineering and software design, are the least explored in the context of LLM applications.
Di Sipio et al. [13] highlighted the lack of an SE classification of PTMs in HF, as the existing one is specific for ML. To address this gap, they defined six macro-tasks and proposed a semi-automated approach to extract SE-relevant information from model cards. However, their validation was limited to just three PTMs (BERT, RoBERTa, and T5), and their categorization did not align with the SDLC.
Yang et al. [17] analyzed the ecosystem of LLMs as of August 2023, curating 366 PTMs from HF for SE. The study also explores the use of LLMs to assist in constructing and analyzing the ecosystem, which increased the model size by 16.5. They focus on code-based LLMs, which represent a more specific scope compared to LLMs for broader SE tasks.
A recent study on benchmarks for code LLMs and agents [18] highlights an imbalance in their distribution across the SDLC. Around 60% of existing benchmarks focus on the software implementation phase, whereas requirements engineering and software design phases are underrepresented.
Despite the notable advances mentioned above, a comprehensive SE-centric classification framework of PTMs in HF remains absent. Existing efforts, such as those by Hou et al. [16] and Wang et al. [18], are grounded in static academic literature and are not designed to analyze or update classifications based on the contents of repositories. Other repository-based studies [13, 17] do not align their categorization with the SDLC or focus narrowly on a subset of code-centric LLMs. Di Sipio et al. [13] take a step toward automation with a semi-automated mapping based on model cards, but their classification remains limited in scope and adaptability. In contrast, this paper adopts an SE perspective to extend the existing taxonomy in Hou et al. [16], to encompass a broader and more diverse set of PTMs, cataloguing them according to specific SE activities and tasks within the SDLC. By analyzing a substantially larger and broader set of PTMs, this work addresses unique SE-specific requirements not met by previous studies. As a result, we propose a novel and exhaustive classification framework tailored to the community.
III Methodology
III-A Research Objectives
Following the Goal Question Metric (GQM) template [20], our goal is to analyze PTMs for the purpose of their classification with respect to their application to SE activities and tasks from the point of view of software engineers in the context of HF.
This goal is structured around two RQs. The first addresses key challenges in automatic classification of PTMs, as this information is essential for the subsequent RQ:
  • RQ1: What challenges arise when automatically classifying SE PTMs in HF?
    - RQ1.1: What is the current documentation status of HF SE PTMs?
    - RQ1.2: What are the key findings regarding the correct classification of SE PTMs?
    - RQ1.3: How stable is this classification information over time?
Next, we examine the role of PTMs in SE, focusing on their applications across different SE activities and tasks:
  • RQ2: How are PTMs utilized for automating SE tasks?
    - RQ2.1: What SE activities and tasks are covered by PTMs?
    - RQ2.2: How are ML tasks related to SE activities?
    - RQ2.3: How are cross-activity PTMs applied across multiple SE activities?
    - RQ2.4: Which base models have been most frequently utilized for SE over time?
    - RQ2.5: What attributes characterize SE PTM in HF?
III-B Research Method
Figure 1: Classification pipeline of PTMs relevant to SE in HF.
We conducted a repository mining study [19] following the pipeline illustrated in Figure 1 to classify PTMs hosted on HF. The subsequent paragraphs detail each step of this process. This approach supports reproducibility, allowing anyone with access to the replication package [21] to validate the results. Additionally, the pipeline is designed to be easily updated, so researchers can reapply it to classify new PTMs added to the platform using the same method. The inclusion criteria are summarized in Table II.
TABLE II: Inclusion criteria to identify SE PTMs in HF.
(1) PTMs must be available in HF.
(2) PTMs must have a valid card description, metadata, or an abstract
from an associated paper on arXiv.
(3) The card description, metadata, or abstract must specify a SE task.
(4) Gemini 2.0 Flash must confirm relevance to an SE activity.
III-B1 Task Identification
To identify SE tasks applicable to the five stages of the SDLC, we built on prior literature and extended it to form a broader taxonomy. We based our approach on Hou et al. [16], which identifies 88 SE tasks. As a preliminary adjustment, we excluded the task agile story point estimation from the software development activity, and we renamed this activity to software implementation for consistency with other studies, such as [11].
Building upon this foundation, we extended the original list of tasks along two dimensions. First, we broadened the scope of SE activities by refining and expanding their definitions. For example, we explicitly considered software architecture tasks as part of the software design activity. Second, we enriched the set of tasks within each activity by incorporating additional tasks drawn from the Software Engineering Body of Knowledge (SWEBOK) [22], resulting in a total of 147 tasks. Specifically, we used the following SWEBOK chapters: requirements engineering (Chapter 1), software design (Chapters 2 and 3), software implementation (Chapter 4), software quality assurance (Chapters 5, 12, and 13), and software maintenance (Chapter 7).
III-B2 Data Collection
We accessed the HF API [23] to retrieve all PTMs available as of March 24, 2025. For each identifier, we tried to retrieve (i) its model card description, a markdown file documenting the model’s characteristics, intended uses, and evaluation [24]; (ii) the associated card metadata, specified in a YAML block (e.g., license, tags, language) [25]; and (iii) the abstract of a linked arXiv paper [26], taking advantage of the cross-platform linking between HF and arXiv [27]. Any of these components may be missing for a given PTM. At that time, there were 1,533,973 PTMs.
III-B3 Data Preparation
To ensure consistency in downstream analysis, we preprocessed the identified tasks and the collected data by applying tokenization, lowercasing, and lemmatization. Tokenization splits the text into smaller units, such as words, which simplifies the subsequent analysis, while lowercasing eliminates inconsistencies arising from variations in capitalization. Additionally, lemmatization reduces words to their base forms, allowing different inflections to be treated as a single entity. Together, these steps standardize the textual descriptions, ensuring greater accuracy in classification.
III-B4 Data Search
As we focused on SE-relevant resources, we searched for the name of each of SE task within the prepared data, resulting in 113,558 PTMs containing at least one of them. When a task consisted of multiple words (e.g., “code generation”), we required all tokens to appear together and in the correct order to be considered a valid match. Moreover, since the search was strictly token-based, task names embedded within other tokens were systematically excluded.
III-B5 Data Filtering
We applied several techniques to refine the available resources and ensure that only the most relevant and accurate ones were included in our classification.
5.1) Outlier Detection: We examined the number of PTMs associated with each task to identify those that significantly deviated from the others. We found that debugging, logging, and coding had notably higher occurrences, together accounting for 89.01% of task occurrences. For debugging, some entries contained debug:None in the metadata, which triggered a match in the search, though it should not have. Similarly, for logging, the lemma “log” often appeared in code snippets or reporting tables, such as those involving logging functions (e.g., import logging) or entries like train log, which were not relevant to the task. For coding, we considered a model to be for coding only if it included the tag “code” in the metadata. To address this, we revisited the raw data for these PTMs without lemmatizing, as we wanted to understand the context in which the lemmas appeared. We checked for any occurrences that did not meet our exclusion criteria (Table III), resulting in a refined set of 14,962 PTMs.
TABLE III: Exclusion criteria for the outlier detection filtering.
Activity Task Description
Software Debugging (i) ‘debug’ after ‘.’, ‘-’, ‘_’, ‘/’, ‘<’, ‘>’
Maintenance (ii) ‘debug’ before ‘-’, ‘_’, ‘/’, ‘:’, ‘<’, ‘>’
(iii) ‘debug’ followed by “ or ’
(iv) “ or ’ followed by ‘debug’
Software Logging (i) ‘log’ followed by an alphabetic
Maintenance character, but not ‘g’
(ii) ‘log’ between a period and a letter
(iii) ‘log’ after ‘.’, ‘-’, ‘_’, ‘/’, ‘(’, ‘)’
(iv) ‘log’ before ‘-’, ‘_’, ‘/’, ‘:’, ‘(’, ‘)’
(v) “import logging”
(vi) “no log”
(vii) “training log”
(viii) “train log”
(ix) “testing log”
(x) “test log”
Software Coding (i) absence of the ‘code’ tag
Implementation
5.2) Unique Card Detection: During preliminary analysis, we observed that many model cards exhibit high textual similarity. For instance, the model card of deepseek-ai/DeepSeek-V3-Base[28] is nearly identical to those of other PTMs on the platform, such as deepseek-ai/DeepSeek-V3[29] and Conexis/DeepSeek-R1-NextN-Channel-INT8 [30]. To address this observation, we applied a duplicate detection step using term frequency - inverse document frequency and cosine similarity. We set a 0.99 similarity threshold to capture near-identical cards while allowing for minimal variations, such as the model name. This resulted in 7,797 unique PTMs, with the full list and their high-similarity matches available in the replication package [21].
5.3) Prompt Pilot Test: As part of the filtering stage, we employed an LLM (Gemini 2.0 Flash, selected due to its strong benchmark performance [31]) to assess the relevance of PTMs to SE activities following the PRIMES framework proposed by De Martino et al. [32]. We conducted five pilot tests (one per SE activity) to validate the LLM. For this reason, we selected five random subsets and ensured representation across all their tasks. The sample sizes were determined using a sample size calculator, applying a 95% confidence level and a 5% margin of error to ensure statistical validity[33]. Table IV presents the unique PTMs per SE activity and the corresponding pilot sample sizes. Note that a given PTM can be associated with more than one SE activity.
TABLE IV: PTMs per SE activity and pilot sample sizes.
SE Activity Unique PTMs Pilot Sample
Requirements Engineering 255 154
Software Design 415 200
Software Implementation 2,993 341
Software Quality Assurance 1,512 307
Software Maintenance 3,255 344
The first author manually annotated all the pilot samples for each SE activity. In addition to providing a binary judgment (yes/no) on whether a PTM was relevant to a given SE activity, the annotator also documented the underlying reasoning behind the decision. To ensure reliability, 10% of the pilot samples were independently annotated by the second and last authors, both knowledgeable in the domain, drawing on established intercoder reliability practices [34]. To further consolidate our annotation guidelines and understanding, the three annotators participated in a joint discussion session focused on the cases where disagreement had occurred. This discussion revealed that some PTMs are inherently challenging to classify due to their broad or ambiguous descriptions. For example, mradermacher/Regex-AI-Llama-3.2-1B-GGUF[35] was initially assigned to software implementation based on the “code generation” tag included by the author in the card metadata. While the description primarily refers to RegEx generation, this tag was an explicit link to the software implementation activity, and we ultimately retained the label. For the 10% of the pilot samples, we applied majority voting to resolve any discrepancies. The resulting decisions, combined with the initial annotations from the first author, are referred to as human annotations.
We then prompted the LLM to provide both a binary judgment and a rationale for each PTM in the pilot set. We used the default temperature setting of the model to balance determinism and variability in responses, ensuring consistent yet informative outputs. The model’s performance was evaluated by comparing its responses with the human annotations using Cohen’s kappa [36]. This process was repeated iteratively, refining the prompt until we reached almost perfect agreement, as seen in Table V. This process was conducted for each of the five SE activities.
TABLE V: Cohen’s kappa for Humans vs LLM agreements.
SE Activity Cohen’s Kappa Agreement
Requirements Engineering 1.00 Perfect
Software Design 0.85 Almost perfect
Software Implementation 0.85 Almost perfect
Software Quality Assurance 0.81 Almost perfect
Software Maintenance 0.82 Almost perfect
Lastly, we applied the five prompts to all resources, resulting in 2,205 unique PTMs meeting our inclusion criteria, ensuring their relevance to SE.
5.4) LLM Generalization Test: To evaluate the LLM’s ability to generalize beyond the initial piloting subset and avoid prompt overfitting, we evaluated it on a separate, manually labeled evaluation subset that was not involved in any piloting activities. The full subset was annotated by the first author, and 10% of this subset was independently labeled. We then computed Cohen’s kappa to measure agreement between the human labels and the LLM’s responses on this subset, observing an almost perfect level of agreement ($k=0.88$).
IV Results
In this section, we present the findings for each of the RQs outlined in Section III-A.
IV-A What challenges arise when automatically classifying SE PTMs in HF? (RQ1)
IV-A1 What is the current documentation status of HF SE PTMs?
As summarized in Table VI, 68.52% of the PTMs have a non-empty model card, and particularly 7.32% mention an SE task. This aligns with prior findings that model information can be missing or inaccurate due to the self-reporting nature of model metrics [37].
To complement this, we analyzed the abstracts of the associated arXiv papers linked to all PTMs. This source revealed that only 0.15% of all PTMs have an associated abstract that mentions SE tasks, while the remaining 99.85% either lack an arXiv ID or do not include mentions to any SE task.
By combining all sources, we identified a total of 113,558 unique PTMs that could be associated with SE tasks. This total includes 1,299 additional PTMs identified exclusively through arXiv abstracts, which were not flagged as SE-relevant based on their model card description and metadata alone.
TABLE VI: Current status of HF PTMs: cards and abstracts.
Source Category PTMs Proportion
Model Card Not available 478,125 31.17%
Available but empty 4,774 0.31%
Available, no SE tasks 938,815 61.20%
Available, with SE tasks 112,259 7.32%
arXiv Not available or no SE tasks 1,531,643 99.85%
abstract Available, with SE tasks 2,330 0.15%
Total unique PTMs with SE tasks 113,558 7.40%
Finding 1.1: The current state of HF shows that SE PTMs scarcely report SE tasks in model cards. Finding 1.2: arXiv abstracts associated with PTMs can provide valuable insights into their SE applicability.
IV-A2 What are the key findings regarding the correct classification of SE PTMs?
The application of our classification pipeline revealed that a significant number of model cards are duplicated. This is primarily due to two reasons: either they correspond to different variants of the same model (e.g., optimized or finetuned versions), or users re-upload PTMs under their own namespace without making significant modifications to the card content. These duplications reflect the existence of an ecosystem in which PTMs depend on and interact with each other [17]. For example, we observed that sparsh35/Meta-Llama-3.1-8B-Instruct [38] has 290 highly similar counterparts, such as triplee/torchtune_8B_full_finetuned_llama3.
1_millfield_241223_meta_before_user_epoch2[39]
. Similarly, the PTM dducnguyen/newdream-sdxl-
20[40]
is associated with 715 near-duplicate PTMs. To mitigate this, we implemented a duplicate detection strategy whereby the earliest published model card is treated as the original, and all others with high textual similarity are labeled as similar variants of that original.
In addition, during the pilot manual annotation phase, we detected that some tasks can be interpreted in different contexts. These ambiguities were addressed using the LLM to ensure consistency in the classification process. Specifically, the LLM was prompted with explicit inclusion and exclusion criteria for each SE activity, allowing it to disambiguate cases based on contextual clues. Notable examples include:
  • Requirements engineering: Terms such as system/software requirement often referred to the technical prerequisites for using the PTM rather than functional requirements in an SE context. Similarly, use case generation was frequently misidentified due to a recurring pattern in many model cards (particularly in Gemma models) where the phrase appeared as part of the “Risks identified and mitigations” section, without referring to requirements engineering.
  • Software design: Expressions like rapid prototyping, architectural design (used in the context of interior design or neural network architectures), and software/system design (commonly present in phrases such as “software designed to…”) did not reflect SE-specific practices but triggered false positives identifications.
  • Software implementation: Many PTMs labeled with API inference or data analysis were unrelated to SE, as they were often linked to generative tasks like Stable Diffusion or generic ML applications.
  • Software quality assurance: The term verification appeared in contexts such as speaker verification, fact verification, or self-verification, which are unrelated to software verification. Similarly, cybersecurity was often mentioned in relation to red teaming assessments.
  • Software maintenance: We found instances where code revision referred to dataset hashes rather than code updates, and sentiment analysis was used in contexts unrelated to software (e.g., general natural language processing tasks).
Finding 1.3: Correct classification of SE PTMs requires addressing duplicated model cards and contextual misinterpretations of SE tasks.
IV-A3 How stable is this classification information over time?
Figure 2 shows a growth in the creation of SE PTMs since 2020, with quarterly data revealing periods of accelerated development. Notably, the number of new PTMs developed for software implementation rose from 28 in 2023 Q1 to 142 in 2023 Q2. This upward trend intensified during the last year (from 2024 Q1 to 2024 Q4), peaking at 444 new PTMs in 2024 Q2. In comparison, the current number of PTMs related to software implementation is more than 20 times higher than that of the next most represented SE activity, software maintenance. In parallel, we observe a steady rise in the number of PTMs associated with requirements engineering, software maintenance, software quality assurance, and software design, indicating a broader diversification of SE tasks being addressed.
Figure 2: Number of SE PTMs created quarterly since 2020.
Finding 1.4: Since 2020, SE PTMs have grown steadily, with sharp acceleration observed from 2023 onward. Finding 1.5: The number of PTMs addressing SE activities continues to grow, with software implementation maintaining a strong lead.
IV-B How are PTMs utilized for automating SE tasks (RQ2)
IV-B1 What SE activities and tasks are covered by PTMs?
Figure 3: PTMs associated with each SE task and SE activity.
Figure 3 shows the distribution of PTMs across SE tasks, with each bar color-coded to reflect its corresponding SE activity. A clear pattern emerges: the most common SE tasks are code generation and coding. Notably, PTMs designed for code generation outnumber those for general coding support by more than a twofold margin. From an activity-level perspective, the data reveal that software implementation is the most dominant SE activity in terms of PTM coverage. In contrast, early-stage activities such as requirements engineering and software design are less represented, with only a small fraction of PTMs targeting them. Later phases of the SDLC, including software quality assurance and maintenance, demonstrate moderate PTM adoption. Tasks like cybersecurity and sentiment analysis are relatively well-supported, indicating a recognition of the value of automation in ensuring software security.
Finding 2.1: Code generation is the most covered SE task. Finding 2.2: Software implementation-related PTMs dominate, while early stages of the SDLC are underrepresented.
IV-B2 How are ML tasks related to SE activities?
The relationship between SE activities and ML tasks is illustrated in Figure 4, with the flow indicating the number of PTMs. The figure highlights the prominence of the ML task text generation, which is associated with the highest number of PTMs, reflecting its relevance in supporting code-related tasks. From the SE perspective, this task is most commonly linked to software implementation. However, it also extends across the remaining four SE activities, indicating its versatility.
Figure 4: Association of SE activities with ML tasks.
ML support is unevenly distributed across the SDLC. Early phases (requirements engineering and software design) and late phases (software maintenance) demonstrate connections to diverse ML tasks, while middle phases like software design show limited ML task integration, being connected primarily to text generation. In particular, software design is associated with a single ML task (text generation) while the other SE activities are supported by multiple ML tasks. Software quality assurance presents an interesting case with connections to several ML tasks, but notably fewer than software implementation or software maintenance. This highlights a potential growth area for ML applications in testing and validation tasks. Regarding software maintenance, it spans across various ML tasks, including text generation, text classification, and text2text generation. These ML tasks are useful in the context of SE, such as code review tasks.
Finding 2.3: The ML task text generation is the most common among SE-related PTMs. Finding 2.4: Unlike other SE activities, software design is related to a single ML task: text generation.
IV-B3 How are cross-activity PTMs applied across multiple SE activities?
PTMs can be related to more than one SE task, potentially spanning different SE activities (i.e., a PTM can be simultaneously classified into multiple SE activities).
To better understand this phenomenon, we selected the PTMs associated with more than one distinct SE activity and represented their pair-wise relationships in Figure 5. The strongest co-occurrence is observed between software implementation and software maintenance, followed by links with software quality assurance and software design. While these numbers highlight absolute overlaps, it is also notable that a substantial proportion of PTMs are reused across activities. For example, 73% of software design PTMs are also applied in software implementation. This suggests that certain PTMs are particularly well-suited for reuse across mid- and late-stage activities in the SDLC, whereas connections to requirements engineering remain relatively sparse.
Figure 5: Co-occurrence graph of cross-activity PTMs.
Finding 2.5: A significant number of PTMs are reused across SE activities, indicating broad applicability. Finding 2.6: Software design PTMs show a particularly strong overlap with software implementation.
IV-B4 Which base models have been most frequently utilized for SE over time?
We analyze the number of SE PTMs developed from specific base models over time using the information provided in the model cards [41]. The “base model” attribute was introduced by HF in November 2023, allowing users to include it in the metadata since then. In this context, users can report whether their PTM is a fine-tuned or adapted version of an original PTM.
Figure 6 presents the top 15 base models with the highest number of PTMs created from them. Established models such as meta-llama/Llama-2-7b-hf[42] have supported 9 PTMs since their introduction, maintaining steady reuse since late 2023. Meanwhile, newer models such as Spestly/Atlas-Flash-7B-Preview[43] have rapidly gained traction, each serving as the base for 36 PTMs within a few months of release. This reflects both the sustained value of proven models and the quick adoption of emerging architectures by the SE community.
Figure 6: Top 15 base models by the number of SE PTMs created.
Finding 2.7: Both established and new base models are widely reused to develop SE PTMs.
IV-B5 What attributes characterize SE PTMs in HF?
SE PTMs can be characterized by a set of core attributes that provide insight into their composition, usage conditions, and performance. To examine how they are documented in practice, we conducted an in-depth analysis of the information provided in the model cards [44] of SE PTMs. Our analysis focused on three main attributes: training dataset, license type, and benchmarks.
Training datasets
High-quality data is essential for effectively training PTMs on domain-specific tasks [45]. To understand how such resources are documented in SE PTMs, we extracted the training dataset information from each model card when available. For the purpose of analysis, we identified the top three training datasets per SE activity based on the number of PTMs that reference each dataset. Notably, if any of these top datasets are reused across multiple SE activities, we explicitly indicate this reuse in Figure 7.
The results reveal clear patterns of dataset reuse across SE tasks. For example, codeparrot/github-code-clean[46] is among the most commonly reused datasets, appearing in 24 PTMs for software implementation and 8 for software quality assurance. This reuse suggests that datasets are not always uniquely tailored to specific SE activities.
In contrast, some datasets show activity-specific use. For instance, jondurbin/airoboros-2.2 [47] appears exclusively in 29 PTMs targeting software implementation, while Tesslate/UIGEN-T1.5-Dataset[48] is used solely in 3 design-related PTMs. Such patterns may reflect more deliberate curation or clearer alignment between dataset content and task requirements.
Despite this, significant gaps remain. In particular, no PTMs associated with requirements engineering specify a training dataset in their model card metadata. This absence may reflect the limited availability of standardized datasets for requirements engineering tasks or a lack of consistent reporting practices for this activity.
Figure 7: Top 3 training datasets per SE activity by number of PTMs trained on them.
Finding 2.8: Training datasets are often reused across multiple SE activities. Finding 2.9: No PTMs targeting requirements engineering report associated training datasets.
Licenses
Licensing considerations have become increasingly important in the development and deployment of PTMs, with calls to incorporate license awareness to prevent copyright risks and unintended reuse [49]. In this context, license information provides insights into the legal and practical usage conditions of each PTM. As illustrated in Figure 8, a wide variety of licenses are applied, reflecting different levels of openness, reuse permissions, and restrictions tailored to diverse needs.
Figure 8: License distribution across SE activities.
Permissive licenses such as Apache 2.0 and MIT dominate, particularly in the software implementation phase, where Apache 2.0 licenses cover a striking 857 PTMs and MIT licenses cover 278 PTMs. This dominance aligns with prior works on HF, underscoring the preference for licenses that enable broad use, modification, and redistribution with minimal barriers [50].
More restrictive licenses, such as the Affero General Public License (AGPL-3.0) and the GNU General Public License (GPL-3.0), appear less frequently but are still notable in specific activities (AGPL-3.0 predominantly in software implementation, and GPL-3.0 also appearing in implementation and maintenance phases).
The presence of licenses such as BSD variants (BSD-2-Clause, BSD-3-Clause) and Creative Commons licenses (e.g., CC-BY-4.0, CC-BY-NC-4.0) illustrates additional options that balance openness with certain usage conditions, like attribution or non-commercial clauses. Overall, this diverse and skewed license distribution reflects a pragmatic balance between openness and control that shapes the adoption and reuse of SE PTMs in the HF ecosystem.
Finding 2.10: Permissive licenses, especially Apache 2.0 and MIT, dominate, facilitating broad adoption and reuse. Finding 2.11: Restrictive licenses are less frequent, but notably present in specific activities.
Benchmarks
PTMs are mastering new benchmarks faster than ever, highlighting the increasing pace of progress in the field [51]. This progress has also reinforced the call for benchmarks that reflect the complexity of real-world development scenarios [52]. Recent surveys of benchmarks for LLMs in SE emphasize the diversity of tasks covered and the need for more standardized evaluation methods [53]. However, there is currently no standardized tag or metadata format in HF that enables consistent reporting of benchmarks across PTMs for straightforward comparison.
Figure 9: Top 20 benchmarks most commonly reported in software implementation PTMs in HF.
After manually reviewing many PTMs during the validation stage (Section III-B5), we decided to automatically search for benchmark tables in model cards. Only PTMs related to software implementation reported benchmarks in tabular form. From these, we identified the top 20 most common benchmarks (Figure 9). The most frequent benchmarks include HumanEval [54] (38 occurrences), Massive Multitask Language Understanding (MMLU) [55] (31), and HellaSwag [56] (31). The lack of a unified standard for reporting benchmarks complicates objective comparisons between PTMs.
Finding 2.12: The absence of standardized benchmark reporting hinders consistent comparison across SE PTMs. Finding 2.13: HumanEval, MMLU, and HellaSwag are the most frequently reported benchmarks among SE PTMs.
V Discussion
A significant gap exists in the documentation of SE PTMs in HF. At the time of the data collection, over 31% of the available PTMs lacked a model card, and only 7.32% included a valid model card that explicitly referenced an SE task. This abscence of documentation reduces transparency, limits informed reuse, and creates obstacles to evaluating the suitability of PTMs for specific SE applications.
Our results align with Hou et al. [16], both identifying a strong focus on code generation tasks, reflecting the broad adoption of generative PTMs in development workflows. As shown in our findings, the number of PTMs associated with software implementation is more than 20 times higher than that of the next most represented SE activity, software maintenance. Our study advances the field by analyzing a substantially larger and more diverse set of over 2,205 PTMs, compared to Hou et al.’s literature-based analysis [16].
Consistent with Yang et al. [17], who focus on code-centric LLMs, we observe this skew toward software implementation. Yet, our work broadens the scope by covering a wider range of SE tasks and phases (identifying 147 distinct SE tasks across five SDLC activities), providing finer granularity and alignment with the full lifecycle.
This imbalance highlights the urgent need for greater research and tooling support for early SDLC activities, such as requirements engineering and software design. The growing interest in PTMs simulating diverse SE roles, such as MetaGPT [57], and the lack of benchmarks for these phases [18], highlight this gap.
Our findings motivate future efforts to extend classification frameworks to additional repositories and develop tools to foster more balanced PTM support across the SDLC, unlocking PTMs’ full potential beyond software implementation.
V-A Implications
The findings of this study provide a practical foundation for both researchers and practitioners seeking to navigate the rapidly growing landscape of SE PTMs. Consider an SE team at a US-based company developing an intelligent code completion tool for Python, integrated into their existing desktop IDE. This team aims to enhance developer productivity through automation while prioritizing environmental sustainability and compliance with open-source licensing policies.
Our classification supports in identifying and selecting SE PTMs that best align with specific needs, thereby streamlining decision-making processes [58]. Using our catalogue, the team begins by filtering available PTMs for those applicable to software implementation, and they retain only those with explicit support for Python. The next filter selects PTMs that report energy consumption during inference, aligning with their sustainable AI practices. From this subset, they identify PTMs released under permissive licenses, such as MIT. When multiple candidates meet all criteria, the team evaluates them based on quality attributes such as accuracy and benchmark performance. Through this process, the team identifies ashwinR/CodeExplainer[59] as the most suitable, demostrating how our framework can support decision-making in real-world scenarios.
Figure 10 depicts this selection process, emphasizing how our approach guides users to find PTMs best suited to their specific context. Ultimately, this example highlights the framework’s potential to streamline the adoption of SE PTMs, promoting responsible AI development aligned with broader organizational and societal goals.
Figure 10: Selection diagram.
V-B Updatability
The replication package [21] accompanying this work includes the complete pipeline described in Section III, enabling full replicability of our results. By following the documented steps, users can reproduce the same classification of PTMs available up to March 24, 2025, while also retrieving and classifying PTMs added since that date. This supports replicability and allows others to validate or extend our work.
However, given the growing number of PTMs in HF [60, 8], it is not necessary to rerun the full pipeline unless modifications are required (e.g., updating the taxonomy of SE tasks or validating the classification strategy). To keep the dataset up to date, users can retrieve only the new PTMs added since our cutoff date and apply the pipeline exclusively to those. The resulting classified PTMs can then be merged with our dataset, allowing efficient updatability.
VI Threats to Validity
Potential threats impacting our study’s validity are outlined.
Internal validity: Can be compromised by the quality of the model card descriptions, as incomplete or ambiguous documentation could lead to misclassification. To mitigate this, we enriched our data collection process by incorporating additional sources of evidence, such as metadata fields and abstracts from linked arXiv publications, thereby improving the robustness of our classification decisions.
External Validity: While HF is a leading platform for sharing PTMs, other repositories, such as PyTorch Hub [61], may host relevant PTMs that are not captured in our analysis. Moreover, ongoing changes in HF’s platform structure, documentation practices, or model metadata could affect the broader applicability of our conclusions.
Construct Validity: Our classification of SE tasks and activities builds upon an established taxonomy derived from existing literature [16]. We extended this taxonomy to include a broader set of 147 SE tasks, aiming for more granular coverage. Nevertheless, it may evolve as the SE community introduces new standards or refinements, which could impact future interpretations or mappings.
Conclusion Validity: We acknowledge the use of Gemini 2.0 Flash to support part of the classification process. To mitigate potential risks, we designed a pilot evaluation protocol with five distinct prompting scenarios (one per SE activity), and validated outputs on a statistically significant sample. A total of 1,346 samples were manually reviewed across all five pilots, involving three independent annotators. This process enabled us to assess the consistency, accuracy, and potential biases in the model’s responses.
To further address these threats, we have made our entire study fully replicable. All data, scripts, and resources are publicly available, allowing future researchers to validate, challenge, or build upon our findings as the ecosystem of SE PTMs continues to evolve.
VII Conclusion and Future Work
This study examines the availability of SE PTMs hosted on HF, identifying 2,205 PTMs relevant to SE. Software design and requirements engineering are under-represented, highlighting a need for broader coverage across the SDLC. The most prevalent ML task across SE-related PTMs is text generation, while code generation is the most commonly addressed SE task. Additionally, there has been a significant surge in SE PTMs since 2023 Q2. This snapshot of the current landscape highlights the existing gaps in SE resources and provides valuable insights into the field’s evolving needs, helping to motivate future research efforts. By classifying PTMs according to SE activities, our work also contributes to automating the sampling and selection of relevant PTMs for researchers and practitioners.
For future work, we plan to extend this classification to other repositories, such as Papers with Code [62], PyTorch Hub [61], and TensorFlow Hub [63], allowing for a broader analysis of PTMs in SE. We also aim to enhance the classification by recovering and integrating relevant information from original base model cards, which may not be explicitly stated in derivative model cards. To make our classification more actionable, we propose developing a dashboard that enables automatic sampling and selection of PTMs for real-world applications, supporting both researchers in conducting empirical studies and practitioners in integrating PTMs into software engineering workflows. Finally, we intend to explore how this classification can be embedded into ASE pipelines by enabling automatic recommendation and adaptation of PTMs to SE tasks, such as design pattern or bug detection.
VIII Artifact availability
To support transparency, reproducibility, and reuse, we provide a publicly available replication package containing all research artifacts used in this study [21]. It includes the taxonomy of 147 software engineering tasks aligned with the SDLC, along with the raw and preprocessed data, and the full pipeline for data collection, preparation, search, and filtering. We also share the manual annotations made by human experts during the study.