### v2

Abstract
Context: Open-source Pre-Trained Models (PTMs) provide extensive resources for various Machine Learning (ML) tasks, yet these resources lack a classification tailored to Software Engineering (SE) needs to support the reliable identification and reuse of models for SE. Objective: To address this gap, we derive a taxonomy encompassing 147 SE tasks and apply an SE-oriented classification to PTMs in a popular open-source ML repository, Hugging Face (HF). Method: Our repository mining study followed a five-phase pipeline: (i) identification SE tasks from the literature; (ii) collection of PTM data from the HF API, including model card descriptions and metadata, and the abstracts of the associated arXiv papers; (iii) text processing to ensure consistency; (iv) a two-phase validation of SE relevance, involving humans and LLM assistance, supported by five pilot studies with human annotators and a generalization test; (v) and data analysis. This process yielded a curated catalogue of 2,205 SE PTMs. Results: We find that most SE PTMs target code generation and coding, emphasizing implementation over early or late development stages. In terms of ML tasks, text generation dominates within SE PTMs. Notably, the number of SE PTMs has increased markedly since 2023 Q2, while evaluation remains limited: only 9.6% report benchmark results, mostly scoring below 50%. Conclusions: Our catalogue reveals documentation and transparency gaps, highlights imbalances across SDLC phases, and provides a foundation for automated SE scenarios, such as the sampling and selection of suitable PTMs.
1 Introduction
The exponential growth of open-source platforms such as Hugging Face (HF) huggingface has enhanced access to Machine Learning (ML) assets zhao2024empirical. In particular, HF is distinguished by its extensive and fast-growing collection of Pre-Trained Models (PTMs) han2021pre; ait2024hfcommunity, compared to other platforms 10.1145/356993410.1145/3643991.3644907. In 2024 alone, HF saw an average of 2,199 new PTMs published per day and over six million daily downloads, based on our analysis of data retrieved through the platform’s public API. Moreover, it is estimated that a new repository is created every 15 seconds  [^1] https://huggingface.co/posts/clem/238420842235482, reflecting trends observed in prior studies that reported rapid increases in repository creation and model contributions 10.1145/3643991.3644898. Despite this growth, the current categorization of HF resources remains largely general-purpose, often targeting ML domains as broad as computer vision or natural language processing. This generic taxonomy fails to address the specific and diverse needs of the Software Engineering (SE) community, whose requirements differ significantly from those of other fields barenkamp2020applications. Importantly, SE goes beyond mere coding, encompassing a set of activities structured around the well-known Software Development Life Cycle (SDLC) ruparelia2010software; sommerville2015. This cycle comprises five main activities: requirements engineering, software design, software implementation, software quality assurance, and software maintenance, thereby covering the entire process from initial specification to long-term system evolution.
For this reason, SE researchers giray2021software and practitioners tan2024challenges; zhao2024empirical face challenges in identifying HF’s ML assets that suit their tasks, while meeting project-specific constraints (e.g., licensing requirements or minimum performance scores). Common issues include missing attributes, discrepancies between reported and actual performance, and risks related to privacy or unethical model behaviour jiang2023empirical. The absence of an SE-specific categorization in HF 10.1145/3661167.3661215 sets a significant barrier, by limiting the efficient application of ML in SE. This gap has severe consequences for SE practitioners and researchers. In the absence of tailored filtering mechanisms, users must manually sift through thousands of PTMs to sample those relevant for their tasks, an effort that is both time-consuming and error-prone, ultimately slowing progress and reducing reproducibility 8804457. Although search engines can facilitate retrieval to some extent, they rely on heterogeneous metadata and lack an SE-specific focus, which favours popular PTMs over those better suited to SE tasks. Moreover, the lack of a task-aligned taxonomy prevents the SE community from systematically identifying relevant resources, evaluating progress, and enabling full automation in the multi-agent environment.
Given the dynamic and evolving nature of both ML and SE 10752650, there is a clear need for a flexible and domain-specific cataloguing system. Such a system would support the effective sampling, selection, and integration of PTMs tailored to the full SDLC. The ability to identify suitable alternatives remains essential for sustainable and adaptable development practices. Moreover, open-source registries such as HF are founded on the principle of choice, offering visibility and reuse opportunities beyond the most popular PTMs. A structured taxonomy can therefore enhance discoverability, guiding users towards PTMs that better fit specific SE needs.
Prior studies have taken important steps toward organizing ML resources for SE, but each presents limitations that leave a significant gap unaddressed. 10.1145/3695988 proposed a taxonomy by categorizing LLM-related papers into SE tasks and SDLC activities through a systematic literature review. However, their approach is grounded in static academic sources and lacks connection to actual open-source registries like HF. 10.1145/3661167.3661215 extracted SE-relevant information from HF model cards using a semi-automated method. Yet, their validation was limited to only three PTMs, and their taxonomy does not align with the SDLC, reducing its applicability. Similarly, yang2024ecosystemlargelanguagemodels curated code-focused LLMs from HF, providing useful insights into that specific subset, but their scope excludes broader SE tasks and lacks an SDLC perspective. citewang2025software surveyed benchmarks for code LLMs and agents around the SDLC. However, their work does not address the classification of other types of PTMs.
In this work, we address these limitations by introducing an automated cataloguing framework for SE tasks across the SDLC. Our approach systematically maps PTMs from HF to a taxonomy of 147 SE tasks, built on prior research and established guidelines. We analyze multiple sources of descriptive content, including model card descriptions, metadata, and abstracts from associated arXiv papers. The pipeline begins with the extraction of candidate task tokens from textual descriptions and applies several processing steps to improve precision. These include outlier detection to remove inconsistent entries, similarity-based deduplication to eliminate near-identical variants, and LLM support to validate that the assigned activity reflects the model’s described purpose. Our main contributions are as follows:
  • (a)
    An automated cataloguing framework that retrieves, processes, and validates PTMs from the HF API, classifying each according to the full SDLC.
  • (b)
    An analysis of SE PTMs, covering their distribution across SE activities and tasks, cross-activity reuse, and their key attributes, including temporal evolution, base models, benchmarks, training datasets, and licenses.
  • (c)
    A validated, publicly available dataset of 2,205 SE PTMs as of March 2025.
Our results offer several key insights. By providing an SDLC-aligned catalogue of PTMs, our work supports researchers and practitioners by automating the identification and selection of the most appropriate PTMs for specific tasks. This automation reduces manual effort, improves consistency in benchmarking, and accelerates the adoption of ML-driven tools within SE workflows, ultimately fostering more effective and targeted automation across the SDLC. In addition, our analysis uncovered that the number of SE PTMs in HF has grown significantly in the past few years, being heavily concentrated on the mid to late stages of the SDLC.
This article is structured as follows. Section 2 reviews related work on PTMs and their application in SE. Section 3 describes the methodology used to catalogue PTMs hosted on HF. Section 4 presents the main results and analysis. Section 5 discusses the findings and explores implications. Section 6 identifies threats to validity; and Section 7 concludes the article and suggests future research directions.
2 Related Work
HF offers a rich ecosystem for conducting empirical studies. Recent repository mining studies have explored the HF ecosystem from several angles to assess its utility and characteristics. ait2025suitability evaluated the suitability and potential of the platform for supporting empirical studies, and jiang2023empirical addressed PTM reuse. Researchers have also investigated standardization and release practices within the ecosystem, including the study of PTM naming conventions by jiang2025see, and the analysis of release transparency, versioning, and naming by ajibode2025towards. Other works have focused on the lifecycle and management of these assets, such as castano2024machine, who conducted a large-scale longitudinal study on how models change over time. While these studies characterize the broader HF landscape, our work specifically targets the identification and cataloguing of SE PTMs.
Key aspects of the related work are summarized in Table 1. A systematic literature review conducted by 10.1145/3695988 analyzed 395 research papers from January 2017 to January 2024 and categorized Large Language Models (LLMs) into SE tasks. These tasks were grouped into SE activities according to the SDLC. The study identified 88 distinct SE tasks, revealing that LLMs are most commonly applied in the software development activity, with code generation and program repair being the most prevalent tasks for employing LLMs in software development and maintenance activities, respectively. In contrast, the earlier SDLC phases, such as requirements engineering and software design, are the least explored in the context of LLM applications.
Recent work by 10.1145/3661167.3661215 highlighted the lack of an SE classification of PTMs in HF, as the existing one is specific for ML. To address this gap, they defined six macro-tasks and proposed a semi-automated approach to extract SE-relevant information from model cards. However, their validation was limited to just three PTMs (BERT, RoBERTa, and T5), and their categorization did not align with the SDLC.
Similarly, yang2024ecosystemlargelanguagemodels analyzed the ecosystem of LLMs as of August 2023, curating 366 PTMs from HF for SE. The study also explores the use of LLMs to assist in constructing and analyzing the ecosystem, which increased the model size by 16.5. They focus on code-based LLMs, which represent a more specific scope compared to LLMs for broader SE tasks.
wang2025software examined benchmarks for code LLMs and agents, revealing an imbalance in their distribution across the SDLC. Around 60% of existing benchmarks focus on the software implementation phase, whereas requirements engineering and software design phases are underrepresented.
Complementary to these efforts, ferino2025novice conducted a systematic literature review of 80 studies to investigate LLM adoption among novice developers. Their findings indicated that LLMs are being used across SDLC phases (except for software management), with most of the addressed tasks belonging to software development and little attention on software design. In their study, they organized the activities according to the SDLC, but grouped some of them together (i.e., requirements engineering and software design, software development and software quality assurance, software maintenance).
niu2022deep provided a survey of PTM for code. They categorized 20 CodePTMs across 18 SE tasks and four dimensions: architecture, modality, training tasks, and programming languages. Their taxonomy emphasized the distinction between understanding and generation tasks, highlighting the expanding role of LLMs in automating software implementation activities.
Table 1: Summary of relevant aspects in related work on cataloguing artificial intelligence models for SE
Study Approach githubEmpiricalStandardsdocsstandardsMaster SDLC Classification Information Used Initial Resources Identified Resources 10.1145/3695988 Systematic Review 88 SE tasks across 6 SE activities Paper content 218,765 papers 395 papers 10.1145/3661167.3661215 Repository Mining 6 macro-tasks HF model card 381,240 PTMs Tested on: BERT, RoBERTa, T5 yang2024ecosystemlargelanguagemodels Repository Mining 9 categories HF model documentation 1,566 PTMs 366 PTMs wang2025software Systematic Review 30 SE tasks across 6 SE activities Academic publications 1,347 papers 181 benchmarks from 461 relevant papers ferino2025novice Systematic Review 35 SE tasks across 5 SE activities Paper content 501 studies 80 studies niu2022deep Survey 18 tasks across understanding and generation Model technical details 20 PTMs 20 PTMs This study Repository Mining 58 SE tasks (from a taxonomy of 147 SE tasks) across 5 SE activities HF model card description, metadata, arXiv abstract 1,533,973 PTMs 2,205 PTMs
Beyond academic studies, community platforms such as paperswithcode also offer structured connections between papers and code implementations, offering valuable categorizations of ML tasks  [^2] In July 2025, Papers With Code was integrated into HF as Trending Papers (https://huggingface.co/papers/trending). Although the original website is no longer available, the historical PWC data remains accessible via HF datasets (e.g., https://huggingface.co/datasets/pwc-archive/papers-with-abstracts).. However, these are primarily oriented toward traditional ML domains. Even though the platform includes an SE-related category (i.e., computer code), it combines general ML tasks (e.g., reinforcement learning, semantic segmentation) with SE-specific tasks (e.g., code generation, program repair), limiting its utility for SE-focused research and practice.
Despite the notable advances mentioned above, a comprehensive SE-centric classification framework of PTMs in HF remains absent. Existing efforts, such as those by 10.1145/3695988 and wang2025software, are grounded in static academic literature and are not designed to analyze or update classifications based on the contents of repositories. Other repository-based studies 10.1145/3661167.3661215; yang2024ecosystemlargelanguagemodels do not align their categorization with the SDLC or focus narrowly on a subset of code-centric LLMs. 10.1145/3661167.3661215 take a step toward automation with a semi-automated mapping based on model cards, but their classification remains limited in scope and adaptability. In contrast, this paper adopts an SE perspective to extend the existing taxonomy in 10.1145/3695988, to encompass a broader and more diverse set of PTMs, cataloguing them according to specific SE activities and tasks within the SDLC. By analyzing a substantially larger and broader set of PTMs, this work addresses unique SE-specific requirements not met by previous studies. As a result, we propose a novel and exhaustive classification framework tailored to the SE community.
3 Methodology
In this section, we define the goal and objectives of our study, together with the research questions that guided our investigation. We also provide the rationale behind our approach and describe the different phases of the study.
3.1 Research Objectives
Following the Goal Question Metric (GQM) template caldiera1994goal, our goal is to analyze HF PTMs for the purpose of their automatic classification with respect to their application to SE activities and tasks from the point of view of software engineers in the context of model selection.
This goal is structured around three RQs. The first focuses on key challenges in the automatic cataloguing of PTMs, providing essential information for the subsequent questions:
  • RQ1: What challenges arise when automatically cataloguing SE PTMs in HF?
    - RQ1.1: What is the current documentation status of SE PTMs?
    - RQ1.2: What are the key findings regarding the correct cataloguing of SE PTMs?
Next, we examine the role of PTMs in SE, focusing on their applications across the entire SDLC:
  • RQ2: How are HF PTMs utilized for supporting SE tasks?
    - RQ2.1: What SE activities and tasks are covered by PTMs?
    - RQ2.2: How are cross-activity PTMs applied across multiple SE activities?
    - RQ2.3: How are ML tasks related to SE activities?
Lastly, we explore selected characteristics of SE PTMs, including temporal evolution, their underlying base model, associated benchmark scores, training datasets, and licensing information, to provide an in-depth understanding of their properties.
  • RQ3: What key attributes characterize SE PTM in HF?
    - RQ3.1: How has the number of SE PTMs evolved over time?
    - RQ3.2: Which base models have been most frequently utilized for SE over time?
    - RQ3.3: What patterns emerge from the benchmark scores of SE PTMs?
    - RQ3.4: Which training datasets are used to develop SE PTMs?
    - RQ3.5: What types of licenses are applied to SE PTMs?
By addressing these RQs, we aim to understand the current landscape of SE PTMs, thereby characterizing the broader SE PTM ecosystem. Furthermore, by building a database that consolidates these PTMs and their metadata, we seek to support SE researchers and practitioners in discovering suitable PTMs for their projects.
3.2 Research Method
Following the ACM/SIGSOFT Empirical Stardards githubEmpiricalStandardsdocsstandardsMaster for mining repositories, we designed a study to classify PTMs hosted on HF using the pipeline illustrated in Figure 1. The process comprises five main phases. In the first phase (Task Identification), we identified 147 SE tasks from the literature to serve as the basis for the cataloging. In the second phase (Data Collection), we extracted information from each PTM available on the HF platform, including its model card description, metadata, and related arXiv abstracts. The third phase (Data Processing) involved several steps to normalize the collected textual data, detect SE task matches, remove outliers, and identify unique model cards to ensure consistency. In the fourth phase (LLM-Based Validation), we used Gemini 2.0 Flash to refine the catalogue by retaining only PTMs closely aligned to SE. Finally, in the fifth phase (Data Analysis), we analyzed the data gathered throughout the pipeline to answer each of the RQs. The following paragraphs describe each phase in detail.
Figure 1: Pipeline used to extract, catalogue and analyze SE PTMs in HF
This method supports reproducibility, allowing anyone with access to the replication package replication_package to validate the results. Additionally, the pipeline is designed to be easily updated, so researchers can reapply it to classify new PTMs added to the platform using the same method. The inclusion criteria for identifying SE PTMs in HF are summarized in Table 2.
Table 2: Inclusion criteria to identify SE PTMs in HF
(1) PTMs must be available in HF.
(2) PTMs must have a valid card description, metadata, or an abstract
from an associated paper on arXiv.
(3) The card description, metadata, or abstract must specify a SE task.
(4) Gemini 2.0 Flash must confirm relevance to an SE activity.
3.2.1 Task Identification
To identify SE tasks applicable to the five stages of the SDLC, we built on prior literature and extended it to form a broader taxonomy. We based our approach on 10.1145/3695988, which identifies 88 SE tasks. As a preliminary adjustment and after discussion among the authors, we excluded the task agile story point estimation from the software development activity, and we renamed this activity to software implementation for consistency with other authors, such as sommerville2015.
Building upon this foundation, we extended the original list of tasks along two dimensions. First, we broadened the scope of SE activities by refining and expanding their definitions. For example, we explicitly considered software architecture tasks as part of the software design activity. Second, we enriched the set of tasks within each activity by incorporating additional tasks drawn from the Software Engineering Body of Knowledge (SWEBOK) washizaki2024swebok. Specifically, we considered key concepts and topics described in relevant chapters: requirements engineering (Chapter 1), software design (Chapters 2 and 3), software implementation (Chapter 4), software quality assurance (Chapters 5, 12, and 13), and software maintenance (Chapter 7). The resulting taxonomy, refined through discussion among the authors to ensure coverage and non-overlap, comprises a total of 147 tasks.
3.2.2 Data Collection
We accessed the HF API huggingfaceHuggingFace to retrieve all PTMs available as of March 24, 2025. For each identifier, we tried to retrieve (i) its model card description, a markdown file documenting the model’s characteristics, intended uses, and evaluation mitchell2019model; (ii) the associated card metadata, specified in a YAML block (e.g., license, tags, language) huggingfaceModelCardsMetadata, which has been shown to help users find appropriate PTMs suryani2025model; and (iii) the abstract of a linked arXiv paper huggingfaceModelCardsPaper, taking advantage of the cross-platform linking between HF and arXiv suryani2024exploration. Any of these components may be missing for a given PTM. At that time, there were 1,533,973 PTMs available in HF.
3.2.3 Data Processing
The third phase of the methodology involves four sequential steps to process the collected data, ranging from text normalization (used to standardize matches during task detection) to identifying outliers and unique instances, ensuring that no PTMs are duplicated or counted twice.
3.1) Text Normalization
We preprocessed the list of SE tasks and the collected data to ensure accurate SE task detection across heterogeneous textual sources. Model cards, metadata, and abstracts in HF exhibit diverse writing styles and formats, as terms may appear in uppercase or inflected forms (e.g., “Code Summarization”, “code summarisation” or “code summarizing”). To handle this variability, we applied tokenization, lowercasing, and lemmatization. Together, these steps standardize the textual descriptions, ensuring greater catalogue accuracy.
3.2) Task Search
As we focused on SE-relevant resources, we searched for the name of each of SE task within the prepared data, resulting in 113,558 PTMs containing at least one of them. When a task consisted of multiple words (e.g., “code generation”), we required all tokens to appear together and in the correct order to be considered a valid match. Moreover, since the search was strictly token-based, task names embedded within other tokens were systematically excluded.
3.3) Outlier Detection
We applied several techniques to refine the available resources and ensure that only the most relevant and accurate ones were included in our catalogue. Our primary focus was on detecting positive outliers, i.e., SE tasks whose frequency was abnormally high compared to the rest, as these cases typically signal false positives caused by lexical ambiguities rather than actual SE relevance. We found that debugging, logging, and coding had notably higher occurrences, together accounting for 89.01% of task occurrences. For debugging, some entries contained debug:None in the metadata, which triggered a match in the search, though it should not have. Similarly, for logging, the lemma “log” often appeared in code snippets or reporting tables, such as those involving logging functions (e.g., import logging) or entries like train log, which were not relevant to the task. For coding, we considered a model to be for coding only if it included the tag “code” in the metadata. To ensure robustness of the filtering, we revisited the raw data for these PTMs without lemmatizing, as we wanted to understand the context in which the lemmas appeared, applying the exclusion criteria in Table 3. We then checked for any occurrences that did not meet our exclusion criteria, resulting in a refined set of 14,962 PTMs.
Table 3: Exclusion criteria for the outlier detection filtering
Activity Task Description
Software Debugging (i) ‘debug’ after ‘.’, ‘-’, ‘_’, ‘/’, ‘<’, ‘>’
Maintenance (ii) ‘debug’ before ‘-’, ‘_’, ‘/’, ‘:’, ‘<’, ‘>’
(iii) ‘debug’ followed by “ or ’
(iv) “ or ’ followed by ‘debug’
Software Logging (i) ‘log’ followed by an alphabetic
Maintenance character, but not ‘g’
(ii) ‘log’ between a period and a letter
(iii) ‘log’ after ‘.’, ‘-’, ‘_’, ‘/’, ‘(’, ‘)’
(iv) ‘log’ before ‘-’, ‘_’, ‘/’, ‘:’, ‘(’, ‘)’
(v) “import logging”
(vi) “no log”
(vii) “training log”
(viii) “train log”
(ix) “testing log”
(x) “test log”
Software Coding (i) absence of the ‘code’ tag
Implementation
3.4) Unique Card Detection
During preliminary analysis, we observed that many model cards exhibit high textual similarity. For instance, the model card of deepseek-ai/DeepSeek-V3-Base huggingfaceDeepseekaiDeepSeekV3BaseHugging is nearly identical to those of other PTMs on the platform, such as deepseek-ai/DeepSeek-V3 huggingfaceDeepseekaiDeepSeekV3Hugging and Conexis/DeepSeek-R1-NextN-Channel-INT8 huggingfaceConexisDeepSeekR1NextNChannelINT8Hugging. In particular, duplicate and near-duplicate PTMs often emerge because users re-upload base models with little or no modification horwitz2025charting. To address this observation, we applied a duplicate detection step using term frequency - inverse document frequency and cosine similarity. We set a 0.99 similarity threshold, which is high enough to detect near-identical cards while allowing true distinct versions to remain separate because their documentation differs sufficiently to fall below this threshold. Without this step, we risk double-counting PTMs that have minimal documentation differences but are otherwise the same model. This procedure resulted in 7,797 unique PTMs, with the full list and their high-similarity matches available in the replication package replication_package.
3.2.4 LLM-Based Validation
In the fourth step of the methodology, we used an LLM to validate the cataloguing and ensure that we only keep SE-relevant resources. To guarantee objectivity, we conducted a two-phase validation procedure, as illustrated in Figure 2. All prompt iterations and manually annotated subsets are available in the replication package replication_package.
Figure 2: Process of using an LLM to validate the catalogue
4.1) Prompts Refinement
We employed Gemini 2.0 Flash (selected due to its strong benchmark performance rahman2025comparative) to assess the relevance of PTMs to SE activities following the PRIMES framework proposed by de2024experiences.
We conducted five pilot tests (one per SE activity) to validate the LLM. For this reason, we selected five random subsets and ensured representation across all their tasks. The sample sizes were determined using a sample size calculator, applying a 95% confidence level and a 5% margin of error to ensure statistical validity qualtricsSampleSize. Table 4 presents the unique PTMs per SE activity and the corresponding pilot sample sizes. Note that a given PTM can be associated with more than one SE activity.
The first author manually annotated all pilot samples for each SE activity. In addition to providing a binary judgment (yes/no) on whether a PTM was relevant to a given SE activity, the annotator also documented the underlying reasoning behind the decision. To ensure reliability, 10% of the pilot samples were independently annotated by the second and fourth authors, both knowledgeable in the domain, drawing on established intercoder reliability (ICR) practices in qualitative research o2020intercoder. To consolidate our annotation guidelines and understanding, the three annotators participated in a joint discussion session focused on the cases where disagreement had occurred. This discussion revealed that some PTMs are inherently challenging to classify due to their broad or ambiguous descriptions. For example, mradermacher/Regex-AI-Llama-3.2-1B-GGUF huggingfaceMradermacherRegexAILlama321BGGUFHugging was initially assigned to software implementation based on the “code generation” tag included by the author in the card metadata. While the description primarily refers to RegEx generation, this tag was an explicit link to the software implementation activity, and we ultimately retained the label. For the 10% of the pilot samples, we applied majority voting to resolve any discrepancies. The resulting decisions, combined with the initial annotations from the first author, are referred to as the ground truth.
Table 4: PTMs per SE activity and pilot sample sizes
SE Activity Unique PTMs Pilot Sample
Requirements Engineering 255 154
Software Design 415 200
Software Implementation 2,993 341
Software Quality Assurance 1,512 307
Software Maintenance 3,255 344
We then prompted the LLM in a zero-shot setting to provide both a binary judgment and a rationale for each PTM in the pilot set. We used the default temperature setting of the model to balance determinism and variability in responses, ensuring consistent yet informative outputs.
Prompt Structure The prompt followed a structured format comprising: 1. Objective: Assess the PTM’s applicability to the target activity. 2. Background context: Provide a brief overview of the SDLC and the specific SE activity being evaluated. 3. Task description: Instructions on how to review and assess the PTM for the target SE activity. 4. Criteria: Explicit rules for determining relevance to the target SE activity. 5. Example output: Format for binary judgment and rationale.
The model’s performance was evaluated by comparing its responses with the human annotations using Cohen’s kappa sim2005kappa. This process was repeated iteratively, refining the prompt until we reached almost perfect agreement, as seen in Table 5. This process was conducted for each of the five SE activities. In total, 5 prompt refinements were required for requirements engineering, 12 for software design, 4 for software implementation, 9 for software quality assurance, and 8 for software maintenance. The refinements mainly focused on providing concrete examples of SE tasks for each activity, explicitly instructing the use of the criteria list when making judgments, and avoiding superficial or ambiguous associations with the activity based on the information provided.
Table 5: Cohen’s kappa for Humans vs LLM agreements
SE Activity Cohen’s Kappa Agreement
Requirements Engineering 1.00 Perfect
Software Design 0.85 Almost perfect
Software Implementation 0.85 Almost perfect
Software Quality Assurance 0.81 Almost perfect
Software Maintenance 0.82 Almost perfect
4.2) Generalization Test
To evaluate the LLM’s ability to generalize beyond the initial piloting subset and avoid prompt overfitting, we evaluated it on a separate, manually labeled evaluation subset that was not involved in any piloting activities. This set consisted of 366 randomly sampled PTMs across all SE activities, a statistically significant size for our population. The ground truth for this subset was created using the same rigorous process as before: the full subset was annotated by the first author, and 10% of this subset was independently labeled. Then, we instructed the LLM with the five refined prompts and computed Cohen’s kappa to measure agreement between human decisions and the LLM’s responses on this subset, observing an almost perfect level of agreement ($k=0.88$). This result demonstrates a high degree of generalizability and provides sufficient confidence to use the LLM to label the remaining 6,085 samples.
Lastly, we applied the five prompts to all resources, resulting in 2,205 unique PTMs meeting our inclusion criteria, ensuring their relevance to SE.
3.2.5 Data Analysis
The final step of the methodology involves analyzing all the data gathered throughout the cataloguing process, from the initial data dump from HF to the final set of 2,205 SE PTMs.
RQ1 Analysis
RQ1 focuses on identifying the main challenges in automatically cataloguing SE PTMs in HF. To address RQ1.1, we relied on the original set of PTMs: the data collected from HF consisting on 1,533,973 PTMs. Within this data, we examined the completeness and quality of its documentation. We computed descriptive statistics on the availability of model cards and associated arXiv abstracts, and whether they included any SE task match.
To study contextual misclassifications (RQ1.2), we analyzed the frequency distribution of SE task tokens across all PTMs and applied outlier detection to isolate abnormally frequent lemmas. The contextual meaning of each outlier was manually reviewed to determine whether occurrences were relevant or false positives. This combination of quantitative and qualitative analysis provided insights into the documentation quality and ambiguity of SE PTM cataloguing.
RQ2 Analysis
For RQ2, we examined the utilization of PTMs across SE activities, the extent to which PTMs span multiple activities, and their alignment with underlying ML tasks. Each PTM was associated with one or more SE tasks from the validated taxonomy of 2,205 PTMs, enabling the characterization of their distribution throughout the SDLC (RQ2.1).
The cross-activity analysis was conducted by examining co-occurrence patterns between activities to identify PTMs associated with multiple ones (RQ2.2). These patterns were represented as a co-occurrence matrix that allowed examination of relationships among activities without assuming exclusivity.
Furthermore, to study the connection between SE activities and ML tasks, we exploited the task-based organization currently used in HF, which categorizes models according to their underlying ML task types, and computed their joint distributions with the SE catalogue (RQ2.3).
RQ3 Analysis
The analyses under RQ3 characterized SE PTMs through temporal, structural, and legal dimensions of SE PTMs. For temporal evolution (RQ3.1), we parsed the creation date field of each PTM, normalized timestamps into quarterly intervals, and computed cumulative growth curves to describe the evolution of SE-related models over time.
To analyze base-model reuse (RQ3.2), we extracted the base_model attribute, introduced by HF in November 2023, and aggregated PTMs by base model and creation date to assess dependency patterns among variants of the same base model (e.g., fine-tuned or derivative models).
For benchmark characterization (RQ3.3), we parsed the model-index section of each card to identify evaluation results. Since naming conventions across model cards are heterogeneous, we applied a rule-based normalization to standardize benchmark and metric names (e.g., unifying benchmark variants such as MMLU and Measuring Massive Multitask Language Understanding, or metrics such as accuracy_norm and normalized_accuracy). This analysis is motivated by the growing interest in SE benchmarks in recent years: for SE-focused LLMs, the number of benchmarks increased from 30 in 2022 to 81 in 2024 hu2025assessing.
To analyze training datasets (RQ3.4), we processed the datasets field in the cards to find patterns in dataset reuse. We mapped dataset usage trends and their alignment with SE activities.
Finally, for licensing (RQ3.5), we extracted license information to assess reuse conditions. Licensing considerations have become increasingly important in the development of PTMs, as recent studies highlight the need for license awareness to mitigate copyright risks lo2023trustworthy. Accordingly, we checked how permissive, copyleft, and restrictive licenses are used per SE activities.
4 Results
In this section, we present the findings from our study, organized according to the three RQs introduced in Section 3.1.
4.1 What challenges arise when automatically cataloguing SE PTMs in HF? (RQ1)
To address RQ1, we examine the main challenges encountered when automatically cataloguing SE PTMs in HF. We first analyze the current documentation status of SE PTMs to assess the completeness and reliability of available model information. Next, we discuss the key findings that emerged during the classification process.
4.1.1 What is the current documentation status of SE PTMs?
As summarized in Table 6, 68.52% of the PTMs have a non-empty model card, and particularly 7.32% mention an SE task. This aligns with prior findings that model information can be missing or inaccurate due to the self-reporting nature of model metrics jiang2023empirical.
Table 6: Availability and content of model cards and arXiv abstracts for PTMs in HF
Source Category PTMs Proportion
Model Card Not available 478,125 31.17%
Available but empty 4,774 0.31%
Available, no SE tasks 938,815 61.20%
Available, with SE tasks 112,259 7.32%
arXiv Abstract Not available or no SE tasks 1,531,643 99.85%
Available, with SE tasks 2,330 0.15%
Total unique PTMs with SE tasks 113,558 7.40%
Total unique PTMs 1,533,973 100.00%
To complement this, we analyzed the abstracts of the associated arXiv papers linked to all PTMs. This source revealed that only 0.15% of all PTMs have an associated abstract that mentions SE tasks, while the remaining 99.85% either lack an arXiv ID or do not include mentions to any SE task.
By combining all sources, we identified a total of 113,558 unique PTMs that could be associated with SE tasks. This total includes 1,299 additional PTMs identified exclusively through arXiv abstracts, which were not flagged as SE-relevant based on their model card description and metadata alone.
Finding 1.1: The current state of HF shows that SE PTMs scarcely report SE tasks in model cards. Finding 1.2: arXiv abstracts associated with PTMs can provide valuable insights into their SE applicability.
4.1.2 What are the key findings regarding the correct cataloguing of SE PTMs?
The application of our classification pipeline revealed that a significant number of model cards are duplicated. After identifying SE task matches and removing outliers, we found that only 52.11% of the PTMs were actually unique based on their documentation. This duplication is primarily due to two reasons: either they correspond to different variants of the same model (e.g., optimized or finetuned versions), or users re-upload PTMs under their own namespace without making significant modifications to the card content. These duplications reflect the existence of an ecosystem in which PTMs depend on and interact with each other yang2024ecosystemlargelanguagemodels. For example, we observed that sparsh35/Meta-Llama-3.1-8B-Instruct huggingfaceSparsh35MetaLlama318BInstructHugging has 290 highly similar counterparts, such as triplee/torchtune_8B_full_finet
uned_llama3.1_millfield_241223_meta_before_user_epoch2
huggingfaceTripleetorchtune_8B_full_finetuned_llama31_millfield_241223_meta_before_user_epoch2Hugging. Similarly, the PTM dducnguyen/newdream-sdxl-20 huggingfaceDducnguyennewdreamsdxl20Hugging is associated with 715 near-duplicate PTMs. To mitigate this, we implemented a duplicate detection strategy whereby the earliest published model card is treated as the original, and all others with high textual similarity are labeled as similar variants of that.
In addition, during the pilot manual annotation phase, we detected that some tasks could be interpreted in different contexts. To ensure consistency, we used the LLM with explicit inclusion and exclusion criteria for each SE activity, allowing it to disambiguate cases based on context. Notable examples include:
  • Requirements engineering: Terms such as system/software requirement often referred to the technical prerequisites for using the PTM rather than functional requirements in an SE context. Similarly, use case generation was frequently misidentified due to a recurring pattern in many model cards (particularly in Gemma models) where the phrase appeared as part of the “Risks identified and mitigations” section, without referring to requirements engineering.
  • Software design: Expressions like rapid prototyping, architectural design (used in the context of interior design or neural network architectures), and software/system design (commonly present in phrases such as “software designed to…”) did not reflect SE-specific practices but triggered false positives identifications.
  • Software implementation: Many PTMs labeled with API inference or data analysis were unrelated to SE, as they were often linked to generative tasks like Stable Diffusion or generic ML applications.
  • Software quality assurance: The term verification appeared in contexts such as speaker verification, fact verification, or self-verification, which are unrelated to software verification. Similarly, cybersecurity was often mentioned in relation to red teaming assessments.
  • Software maintenance: We found instances where code revision referred to dataset hashes rather than code updates, and sentiment analysis was used in contexts unrelated to software (e.g., general natural language processing tasks).
Finding 1.3: Correct classification of SE PTMs requires addressing duplicated model cards and contextual misinterpretations of SE tasks.
4.2 How are HF PTMs utilized for supporting SE tasks (RQ2)
To answer RQ2, we analyze how PTMs are applied across different SE activities and tasks. We first examine the coverage of SE tasks to identify which areas of the SDLC are supported by PTMs. Next, we explore PTMs that span multiple SE activities to understand cross-activity reuse, and finally, we investigate the relationships between SE activities and ML tasks to highlight patterns of ML support across the software lifecycle.
4.2.1 What SE activities and tasks are covered by PTMs?
Our taxonomy included 147 SE tasks (Section 3.2.1) across the five SE activities of the SDLC, but only a subset of them was represented among the PTMs identified in HF. After completing all phases of our pipeline, we found that 58 SE tasks were covered in the final validated set of 2,205 SE PTMs. Figure 3 shows the distribution of PTMs across these tasks, with each bar color-coded to reflect its corresponding SE activity.
A clear pattern emerges: code generation and coding are by far the most represented SE tasks. Notably, PTMs designed for code generation outnumber those for general coding support by more than a twofold margin. This highlights the strong focus of text-to-code and code-to-code applications of PTMs.
From an activity-level perspective, the data reveal that software implementation is the most dominant SE activity in terms of PTM coverage. In contrast, early-stage activities such as requirements engineering and software design are less represented, with only a small fraction of PTMs targeting them. Later phases of the SDLC, including software quality assurance and maintenance, demonstrate moderate PTM adoption, with 150 and 159 PTMs supporting them, respectively. Tasks supported in these phases, such as cybersecurity and sentiment analysis are relatively well-supported, indicating a recognition of the value of automation in ensuring software security.
Figure 3: PTMs associated with each SE task and SE activity
Finding 2.1: Code generation and coding are the most covered SE tasks. Finding 2.2: Software implementation-related PTMs outnumber those for early and late SDLC stages.
4.2.2 How are cross-activity PTMs applied across multiple SE activities?
PTMs can be related to more than one SE task, potentially spanning different SE activities (i.e., a PTM can be simultaneously classified into multiple SE activities). To better understand this phenomenon, we selected the PTMs associated with more than one distinct SE activity and represented their pair-wise relationships in Figure 4. The strongest co-occurrence is observed between software implementation and software maintenance, followed by links with software quality assurance and software design.
Although the absolute number of shared PTMs is small relative to the overall corpus of 2,205 SE models, the overlaps are substantial within certain activities. For instance, 22 out of 37 software design PTMs (73%) are also applied in software implementation. This suggests that certain PTMs are particularly well-suited for reuse across mid- and late-stage activities in the SDLC, whereas connections to requirements engineering remain relatively sparse.
Figure 4: Co-occurrence graph of cross-activity PTMs
Finding 2.3: Shared PTMs are few in absolute terms but represent a substantial proportion within certain activities, such as software design.
4.2.3 How are ML tasks related to SE activities?
The relationship between SE activities and ML tasks is illustrated in Figure 5, with the flow indicating the number of PTMs. ML tasks correspond to the current classification used in HF, i.e., a label assigned to each PTM. The figure highlights the prominence of the ML task text generation, which is associated with the highest number of PTMs, reflecting its relevance in supporting code-related tasks. From the SE perspective, this task is most commonly linked to software implementation. However, it also extends across the remaining four SE activities, indicating its versatility.
ML support is uneven across the SDLC. Early phases (requirements engineering and software design) and late phases (software maintenance) demonstrate connections to diverse ML tasks, while middle phases like software design show limited ML task integration, being connected primarily to text generation. In particular, software design is associated with a single ML task (text generation) while the other SE activities are supported by multiple ML tasks. Software quality assurance presents an interesting case with connections to several ML tasks, but notably fewer than software implementation or software maintenance. This highlights a potential growth area for ML applications in testing and validation tasks. Regarding software maintenance, it spans across various ML tasks, including text generation, text classification, and text2text generation. These ML tasks are useful in the context of SE, such as code review tasks.
Figure 5: Association of SE activities with ML tasks
Finding 2.4: The ML task text generation is the most common among SE-related PTMs. Finding 2.5: Unlike other SE activities, software design is related to a single ML task: text generation.
4.3 What attributes characterize SE PTMs in HF? (RQ3)
SE PTMs can be characterized by a set of attributes that provide insights into their properties, performance, and usage conditions. To examine how these attributes are documented in practice, we conducted an in-depth analysis of the information reported in HF model cards huggingfaceModelCards. Our analysis focused on five main attributes: temporal evolution, base model, benchmark scores, training dataset, and license type.
4.3.1 How has the number of SE PTMs evolved over time?
Figure 6 shows a growth in the creation of SE PTMs since 2020, with quarterly data revealing periods of acceleration. Notably, the number of new PTMs developed for software implementation rose from 28 in 2023 Q1 to 142 in 2023 Q2. This upward trend intensified during the last year (from 2024 Q1 to 2024 Q4), peaking at 444 new PTMs in 2024 Q2. In comparison, the current number of PTMs related to software implementation is more than 20 times higher than that of the next most represented SE activity, software maintenance.
In parallel, we observe a steady rise in the number of PTMs associated with requirements engineering, software maintenance, software quality assurance, and software design. Although the absolute number of PTMs for these activities remains much lower than that of software implementation (as shown in Figure 3), the temporal trends indicate a gradual growth across all these SE tasks. This suggests that, while software implementation dominates in volume, other activities are increasingly being addressed by new PTMs.
Figure 6: Number of SE PTMs created quarterly since 2020
Finding 3.1: Since 2020, SE PTMs have grown steadily, with sharp acceleration observed from 2023 onward. Finding 3.2: The number of PTMs addressing SE activities continues to grow, with software implementation maintaining a strong lead.
4.3.2 Which base models have been most frequently utilized for SE over time?
We analyze the number of SE PTMs developed from specific base models over time using the information provided in the model cards huggingfaceModelCardsBaseModel. Figure 7 presents the top 15 base models with the highest number of PTMs created from them.
Figure 7: Top 15 base models by the number of SE PTMs created
Established models such as meta-llama/Llama-2-7b-hf huggingfaceMetallamaLlama27bhfHugging have supported 9 PTMs since their introduction, maintaining steady reuse since late 2023. Meanwhile, newer models such as Spestly/Atlas-Flash-7B-
Preview
huggingfaceSpestlyAtlasFlash7BPreviewHugging have rapidly gained traction, each serving as the base for 36 PTMs within a few months of release. This reflects both the sustained value of proven models and the quick adoption of emerging architectures by the SE community.
Finding 3.3: Both established and new base models are widely reused to develop SE PTMs.
4.3.3 What patterns emerge from the benchmark scores of SE PTMs?
We identified 211 PTMs that included evaluation information (9.57% of the total SE PTMs). Figure 8 illustrates the scores for the ten most frequently reported benchmarks, covering 94.31% of these evaluated PTMs. The distribution reveals that HumanEval, XNLI, and MMLU dominate SE evaluation, accounting for the highest number of reported results, with 1,024, 105, and 103 records, respectively. Notably, despite the large number of PTMs evaluated, most achieve scores below 50% on benchmarks such as BBH, HumanEval, LiveCodeBench, MMLU, MultiPL-HumanEval, RepoQA, and XNLI, indicating that many tasks remain largely unsolved for current PTMs. Among all benchmarks, LiveCodeBench and MultiPL-HumanEval exhibit comparatively lower performances, indicating the difficulty of precise code generation and multi-language evaluation tasks. Conversely, some PTMs reach high scores, up to 94.37% and 90.00% on benchmarks such as XStoryCloze and XCOPA.
Figure 8: Score distribution for the top ten most frequently reported benchmarks in SE PTMs
Finding 3.4: Only 9.57% of SE PTMs report evaluation results, highlighting limited coverage across the ecosystem. Finding 3.5: Most PTMs score below 50% on several common benchmarks, showing many tasks remain largely unsolved. Finding 3.6: XStoryCloze and XCOPA achieve up to 94.37% and 90.00% scores, showing some benchmarks are more tractable for current PTMs.
4.3.4 Which training datasets are used to develop SE PTMs?
For the purpose of analysis, we identified the top three training datasets per SE activity based on the number of PTMs that reference each dataset. Notably, if any of these top datasets are reused across multiple SE activities, we explicitly indicate this reuse in Figure 9.
Figure 9: Top 3 training datasets per SE activity by number of PTMs trained on them
The results reveal clear patterns of dataset reuse across SE tasks. For example, codeparrot/github-code-clean huggingfaceCodeparrotgithubcodecleanDatasets is among the most commonly reused datasets, appearing in 24 PTMs for software implementation and 8 for software quality assurance. This reuse suggests that datasets are not always uniquely tailored to specific SE activities.
In contrast, some datasets show activity-specific use. For instance, jondurbi
n/airoboros-2.2
huggingfaceJondurbinairoboros22Datasets appears exclusively in 29 PTMs targeting software implementation, while Tesslate/UIGEN-T1.5-Dataset huggingfaceTesslateUIGENT15DatasetDatasets is used solely in 3 design-related PTMs. Such patterns may reflect more deliberate curation or clearer alignment between dataset content and task requirements.
Despite this, significant gaps remain. In particular, no PTMs associated with requirements engineering specify a training dataset in their model card metadata. This absence may reflect the limited availability of standardized datasets for requirements engineering tasks or a lack of consistent reporting practices for this activity.
Finding 3.7: Training datasets are often reused across multiple SE activities. Finding 3.8: No PTMs targeting requirements engineering report associated training datasets.
4.3.5 What types of licenses are applied to SE PTMs?
As illustrated in Figure 10, a wide variety of licenses are applied, reflecting different levels of reuse permissions and restrictions tailored to diverse needs.
Figure 10: License distribution across SE activities
Permissive licenses such as Apache 2.0 and MIT dominate, particularly in the software implementation phase, where Apache 2.0 licenses cover a striking 857 PTMs and MIT licenses cover 278 PTMs. This dominance aligns with prior works on HF, underscoring the preference for licenses that enable broad use, modification, and redistribution with minimal barriers pepe2024hugging.
More restrictive licenses, such as the Affero General Public License (AGPL-3.0) and the GNU General Public License (GPL-3.0), are less common but notable: AGPL-3.0 predominantly in software implementation, and GPL-3.0 in implementation and maintenance phases.
The presence of licenses such as BSD variants (BSD-2-Clause, BSD-3-Clause) and Creative Commons licenses (e.g., CC-BY-4.0, CC-BY-NC-4.0) illustrates additional options that balance openness with certain usage conditions, like attribution or non-commercial clauses. Overall, this diverse and skewed license distribution reflects a pragmatic balance between openness and control that shapes the adoption and reuse of SE PTMs in the HF ecosystem.
Finding 3.9: Permissive licenses, especially Apache 2.0 and MIT, dominate, facilitating broad adoption and reuse. Finding 3.10: Restrictive licenses are less frequent, but notably present in specific activities.
5 Discussion
In this paper, we present a large-scale catalogue of 2,205 SE PTMs available in HF, classified according to a fine-grained list of 147 SE tasks derived from the literature 10.1145/3695988washizaki2024swebok. This classification provides significant insights into SE PTMs, their characteristics, and integration within the SDLC. This section discusses identified documentation and transparency gaps, compares our findings with existing prior studies and large-scale HF datasets, and highlights the practical and research implications of our generated catalogue. Lastly, we detail how our methodology can be updated for future extensions.
5.1 Documentation gaps and transparency in SE PTMs
A significant gap exists in the documentation of SE PTMs in HF. Over 31% of the available PTMs lacked a model card, and only 7.32% included a valid model card that explicitly referenced an SE task. This finding reinforces earlier observations that HF model information is frequently incomplete or inaccurate due to its self-reporting nature of the platform jiang2023empirical. Importantly, this issue is not confined to small or experimental PTMs, but also affects large projects with thousands of downloads stalnaker2025ml.
Prior studies confirm that incomplete documentation is a systemic problem across HF. horwitz2025charting, in their analysis of 63,000 models from HF, found that only 8% of them outlined license information, 13.4% reported accuracy details, and 30.2% specified their base model. Similarly, suryani2025model, in a broader-scale study covering the entire HF, reported that nearly one-third of model cards include licensing information. In our SE-focused dataset, only 9.57% of PTMs report evaluation results, often using different metrics and benchmarks, further complicating comparative assessment. Licensing patterns among SE PTMs also reflect broader trends across the HF platform. As reported by stalnaker2025ml, 62.5% of all HF models are released under open-source licenses, with apache-2.0 and mit emerging as the most common choices, consistent with our findings for SE PTMs. Notably, in our data, the “other” license category ranks as the third most frequent among SE PTMs, while in the overall HF corpus it ranks fifth stalnaker2025ml. As discussed by stalnaker2025ml, this category often appears when PTMs were uploaded before the correct license was included in HF’s official list or when model owners are unfamiliar with available licensing options.
These findings align with ecosystem-level trends identified by laufer2025anatomy, who found that model cards within families tend to shorten and become increasingly standardized through templates or automatically generated text laufer2025anatomy, mirroring our observation of model card missing information and similar documentation across PTMs. This absence of documentation reduces transparency, limits informed reuse, and creates obstacles to evaluating the suitability of PTMs for specific SE applications.
5.2 Comparison with prior studies
Our results align with 10.1145/3695988, both identifying a strong focus on code generation tasks, reflecting the broad adoption of generative PTMs in development workflows. As shown in our findings, the number of PTMs associated with software implementation is more than 20 times higher than that of the next most represented SE activity, software maintenance. Our study advances the field by analyzing a substantially larger and more diverse set of over 2,205 PTMs, compared to the literature-based analysis by 10.1145/3695988.
Consistent with yang2024ecosystemlargelanguagemodels, who focus on code-centric LLMs, we observe this skew toward software implementation. Yet, our work broadens the scope by covering a wider range of SE tasks and phases (identifying 147 distinct SE tasks across five SDLC activities), providing finer granularity and alignment with the full lifecycle. This imbalance highlights the urgent need for greater research and tooling support for early SDLC activities, such as requirements engineering and software design. The growing interest in PTMs simulating diverse SE roles, such as MetaGPT hong2023metagpt, and the lack of benchmarks for these phases wang2025software, highlight this gap.
ferino2025novice focused on LLM adoption among novice developers, highlighting that the most addressed tasks belonged to the software development activity and software maintenance. At the same time, they also identified challenges (i.e., losing learning opportunities) and suggested best practices (e.g., prompt engineering). Nevertheless, their study concentrates only on LLMs, with a focus on novice developers’ practices.
In terms of SDLC representation, Figure 11 illustrates the number of models identified by each of the three aforementioned studies (10.1145/3695988; yang2024ecosystemlargelanguagemodels; ferino2025novice) and ours. All four studies show that software implementation is the most dominant SE activity in terms of model coverage, a trend that is most pronounced in our study, with 2,004 PTMs.
Figure 11: Distribution of models per SE activity among related works that are aligned to the SDLC
Regarding requirements engineering, we have catalogued a similar amount of resources as 10.1145/3695988, but significantly less than ferino2025novice. Notably, ferino2025novice grouped multiple SE tasks into broader SE activities, which, for instance, included 20 studies employing LLMs for benchmarking as part of the combined activity “Requirement Design & Software Design”.
For software design, all previous studies analyzed a similar number of models (4, 5, and 7), while we found up to four times more PTMs in HF for such activity (30 PTMs), which may be explained by our broader scope and the inclusion of more recent PTMs.
For the case of software quality assurance, our study identified a considerable jump, cataloging 129 models, significantly surpassing the next highest count of 99 by 10.1145/3695988, and exceeding the counts of 25 and 24 from wang2025software and ferino2025novice, respectively. This suggests greater activity for software quality assurance that previous studies may have missed.
Regarding software maintenance, our study again registered the highest count, with 148 models, compared to 66 for 10.1145/3695988, 24 for wang2025software, and 73 for ferino2025novice This indicates that, while still far behind implementation, maintenance is a growing area for SE PTMs.
The overwhelming prevalence of software implementation PTMs (2,704 models) compared to early-stage activities like requirements engineering (27 models) warrants specific discussion. This disparity is likely driven by the abundance of structured code datasets and mature execution-based benchmarks which facilitate development.
5.3 Overlap with large HF datasets
We have also compared how our catalogue overlaps with two large, generic HF-based datasets to contextualize its coverage. HFCommunity 10123660 aggregates data from the HF Hub, while PeaTMOSS jiang2024peatmoss comprises metadata for 281,638 PTMs and detailed snapshots for those with over 50 monthly downloads (14,296 PTMs), and links PTMs with 28,575 open-source software repositories from GitHub that use them.
We found that $1,338$ ($43.48\%$) and 382382 ($12.38\%$) of our SE PTMs appear in HFCommunity and PeaTMOSS, respectively. This overlap ($0.2\%$ of those collections), arises from the recency of our catalogue, as HFCommunity has been static since October 2024 and PeaTMOSS since August 2023, and from the inclusivity of our catalogue, which retains specialized and emerging PTMs regardless of their popularity.
Table 7 quantifies this temporal gap by illustrating the number of PTMs in our study released after the cutoff dates of the comparison datasets. As both have been static for a long period, they miss a significant portion of the recent surge in SE models. For instance, we identified 2,332 software implementation models and 142 quality assurance models that were published after the PeaTMOSS snapshot. Even compared to the more recent HFCommunity, our catalogue captures 921 additional implementation models released after October 2024. This demonstrates the rapid evolution of the ecosystem. However, these partial intersections create opportunities for richer analysis, since combining our catalogue with broader datasets can reveal ecosystem-level insights that remain hidden when these resources are considered in isolation.
Table 7: Number of SE PTMs in our catalogue published after the data collection cutoff dates of PeaTMOSS and HFCommunity
SE Activity
After PeaTMOSS
(Aug 2023)
After HFCommunity
(Oct 2024)
Requirements engineering 18 8
Software design 36 24
Software implementation 2,332 921
Software maintenance 130 59
Software quality assurance 142 63
5.4 Implications
The findings of this study provide a practical foundation for both researchers and practitioners seeking to navigate the rapidly growing landscape of SE PTMs. By aligning PTMs with SDLC, our catalogue supports informed selection based on project needs, such as license type or reported evaluation.
From a practitioner’s perspective, our framework facilitates the responsible selection and reuse of PTMs in software projects. Consider an SE team at a US-based company developing a code completion tool for Python, integrated into their existing desktop IDE. This team aims to enhance developer productivity through automation while prioritizing environmental sustainability and compliance with open-source licensing policies. Our catalogue supports selecting SE PTMs that best align with specific needs, thereby streamlining decision-making processes ben2025should. As characterized by zhouunifying, a combination of semantic and metric-based selection criteria can be set. In this case, the team queries the database for PTMs catalogued under software implementation and refines the results to those explicitly supporting Python. They then restrict the selection to PTMs that report energy consumption, aligning with their sustainable Artificial Intelligence (AI) practices. From this subset, they identify PTMs released under permissive licenses, such as MIT. When multiple candidates meet all criteria, the team evaluates them based on quality attributes such as accuracy and benchmark performance. Through this process, the team identifies ashwinR/CodeExplainer huggingfaceAshwinRCodeExplainerHugging as the most suitable, demonstrating how our framework can support decision-making in real-world scenarios. Figure 12 depicts this selection process, emphasizing how our approach guides users to find PTMs best suited to their specific context. Ultimately, this example highlights the framework’s potential to streamline the adoption of SE PTMs, promoting responsible AI development aligned with broader organizational and societal goals.
Figure 12: Selection diagram
The findings of this study have also motivated the ongoing development of MLAssetSelection, a web-application that operationalizes our cataloguing pipeline to assist in the sampling and selection of SE PTMs mlassetselection. As illustrated in Figure 13, the tool provides an interactive interface for exploring the catalogue. Users can filter PTMs by identifier, scope (e.g., SE activities and tasks, ML tasks), metadata (e.g., licenses, libraries, natural languages, creation date, size categories: bytes for models and rows for datasets), popularity (e.g., downloads, likes), activity (e.g., commits, contributors), and model-specific attributes (e.g., training dataset, region, inference providers, evaluation).
Figure 13: MLAssetSelection overview
From a research perspective, our findings reveal important gaps and opportunities in the current SE PTM ecosystem. Certain phases of the SDLC remain underrepresented, particularly requirements engineering and software design, where few PTMs explicitly target these activities. Beyond identifying such gaps, our catalogue provides a foundation for researchers to delve deeper into the challenges uncovered in this study and to conduct empirical studies using the collected SE PTMs.
5.5 Updatability
The replication package replication_package accompanying this work includes the complete pipeline described in Section 3, enabling full replicability of our results. By following the documented steps, users can reproduce the same cataloguing of PTMs available up to March 24, 2025, while also retrieving and classifying PTMs added since that date. This supports replicability and allows others to validate or extend our work.
Given the exponential growth of PTMs in HF ait2025suitability; 10.1145/3643991.3644898, re-executing the full pipeline for the entire repository is computationally inefficient and arguably unnecessary unless the taxonomy itself requires modification (e.g., updating the taxonomy of SE tasks or validating the classification strategy). To maintain an up-to-date dataset, our framework supports incremental updates: users can retrieve only the PTMs created after the cutoff date, catalogue them, and merge the results with our baseline dataset. This ensures efficient maintainability without redundant processing.
We acknowledge that a research paper is inherently static (our analysis captures the state of HF as of March 24, 2025), whereas the HF ecosystem is highly dynamic. Since our data collection, the number of repositories has continued to grow at a rapid pace laufer2025anatomy. Given that hundreds of new repositories are created daily  [^3] https://huggingface.co/posts/clem/238420842235482, the absolute numbers presented in this article will inevitably become outdated (as in any other article of this kind). To address this temporal limitation, the value of this work lies not only in the static snapshot provided but in the updatability of the framework. While the raw numbers presented in this article reflect the state of HF as of March 2025, we argue that the distribution of SE tasks and the utility of our taxonomy remain robust. By applying our validated filtering and cataloguing methodology to new entries, the community can maintain a living catalogue. This ensures that the insights regarding the distribution of SE tasks and the utility of the taxonomy remain relevant, even as the underlying raw numbers of the HF ecosystem continue to rise.
Furthermore, our catalogue is designed to be extensible and interoperable with other existing datasets. Although our work focuses specifically on SE, it can be integrated with broader, generic catalogues such as HFCommunity 10123660 or PeaTMOSS jiang2024peatmoss.
6 Threats to Validity
Potential threats that could impact our study’s validity are outlined.
  • Internal validity: Can be compromised by the quality of the model card descriptions, as incomplete or ambiguous documentation could lead to misclassification. To mitigate this, we enriched our data collection process by incorporating additional sources of evidence, such as metadata fields and abstracts from linked arXiv publications, thereby improving the robustness of our classification decisions.
  • External Validity: While HF is a leading platform for sharing PTMs, other registries, such as PyTorch Hub pytorchPyTorch, may host relevant PTMs that are not captured in our analysis. Moreover, ongoing changes in HF’s platform structure, documentation practices, or model metadata could affect the broader applicability of our conclusions.
  • Construct Validity: Our catalogue of SE tasks and activities builds upon an established taxonomy derived from existing literature 10.1145/3695988. We extended this taxonomy to include a broader set of 147 SE tasks, aiming for more granular coverage. Nevertheless, it may evolve as the SE community introduces new standards or refinements, which could impact future interpretations or mappings.
  • Conclusion Validity: While manual annotation may introduce subjective bias, we mitigated this threat through inter-coder agreement practices. We acknowledge the use of Gemini 2.0 Flash to support part of the classification process. To mitigate potential risks, we designed a pilot evaluation protocol with five distinct prompting scenarios (one per SE activity), and validated outputs on a statistically significant sample. A total of 1,346 samples were manually reviewed across all five pilots, involving three independent annotators. This process enabled us to assess the consistency, accuracy, and potential biases in the model’s responses.
To further address these threats, we have made our entire study fully replicable. All data, scripts, and resources are publicly available, allowing future researchers to validate, challenge, or build upon our findings as the ecosystem of SE PTMs continues to evolve.
7 Conclusion and Future Work
This study examines the availability of SE PTMs hosted on HF, identifying 2,205 PTMs relevant to SE. Software design and requirements engineering are under-represented, highlighting a need for broader coverage across the SDLC. The most prevalent ML task across SE-related PTMs is text generation, while code generation is the most commonly addressed SE task. Additionally, there has been a significant surge in SE PTMs since 2023 Q2. This snapshot of the current landscape highlights the existing gaps in SE resources and provides valuable insights into the field’s evolving needs, helping to motivate future research efforts. By classifying PTMs according to SE activities, our work also contributes to automating the sampling and selection of relevant PTMs for researchers and practitioners.
For future work, we plan to extend this classification to other registries, such as GitHub Models githubGitHubModels and PyTorch Hub pytorchPyTorch, allowing for a broader analysis of SE PTMs. We also aim to enhance the catalogue by recovering and integrating relevant information from original base model cards, which may not be explicitly stated in derivative model cards. Finally, we intend to explore how this classification can be embedded into SE pipelines by enabling automatic recommendation and adaptation of PTMs to SE tasks, such as design pattern or bug detection.
Statements and Declarations
Funding
This work was supported by Grant PID2024-156019OB-I00 funded by MICIU/AEI/10.13039/501100011033 and by ERDF, EU. Alexandra González additionally thanks the FI-STEP grant 2025 STEP-00407.
Ethical approval
This study does not involve human participants or animals.
Informed consent
No human subjects were involved in this study.
Author Contributions
  • Alexandra González: Data Collection, Data Processing, Data Validation, Data Analysis, Writing - Original Draft.
  • Xavier Franch: Supervision, Data Validation, Writing - Review & Editing
  • David Lo: Writing - Review & Editing
  • Silverio Martínez-Fernández: Supervision, Data Validation, Writing - Review & Editing
Data Availability Statement
To support transparency, reproducibility, and replicability, and in line with the EMSE open-science guidelines  [^4] https://github.com/emsejournal/openscience, we provide a publicly available replication package containing all research artifacts used in this study replication_package. It includes the taxonomy of 147 software engineering tasks aligned with the SDLC, along with raw and preprocessed data, and the complete pipeline for data collection, preparation, search, and filtering. We also share the manual annotations made by human experts during the study.
Conflict of Interest
The authors declare that they have no conflict of interest.
Clinical Trial Number
Not applicable.